# bHUN-REN Wigner Research Centre for Physics, Konkoly-Thege Miklos ut 29-33, 1121 Budapest, Hungary cInstitut de Physique Théorique [^1] Unité Mixte de Recherche 3681 du CNRS, Université Paris Saclay, CNRS, 91191 Gif-sur-Yvette, France

### v2

@@ 3.1 Properties of $\psi$-function
This operator possesses a discrete spectrum of real, positive eigenvalues that accumulate at the origin Hatsuda:2012hm. These eigenvalues can be conveniently parameterized as $\lambda_{s}=e^{-\varepsilon_{s}}$, where $\varepsilon_{s}$ are monotonically increasing with $s$.
Using the operator representation (16), one finds that the eigenfunctions satisfy a finite-difference equation (see Marino:2016new)
@@
with the expansion coefficients $c_{s}$ independent on $x$ and $z$. It then follows from the last relation that $\psi(x|z)$ is a meromorphic function of $z$ with poles located at the real positive $z=e^{\varepsilon_{s}}$.
@@
This function is entire in $z$ and has zeros located on the negative real axis, at the positions of the poles of the function $\psi(x|-z)$. For reason that will become clear shortly, it is convenient to introduce the ratio of determinants
@@ 3.3 Wronskian relation
Inspection of the integral representation of the matrix elements on the right-hans side shows that the pole at $x-y=i\pi k$ arises solely from the $O(z)$ term. Matching the residues of (51) and (52) at this pole then yields the relation
@@ 4 Quantization condition
In the previous section, we expressed the generating function (3) in terms of the auxiliary $\psi$-function and derived a set of nontrivial relations satisfied by this function. These relations are exact and hold for arbitrary values of the Cher–Simons level $k$, the winding number $n$, and the fugacity parameter $z$.
@@ 4.3 Phase at odd $k$
$\begin{pmatrix}\psi(i\pi k|z)\\[5.69054pt] \psi(-i\pi k|z)\end{pmatrix}=\mathbb{T}_{\rm o}\begin{pmatrix}\psi(i\pi|z)\\[5.69054pt] \psi(-i\pi|z)\end{pmatrix}$, (78)
where the corresponding transfer matrix is given by
@@ 4.4 Solutions to the quantization conditions
It is straightforward to verify that the last two relations in (4.4) imply the same asymptotic behaviour for $k=3,4$. Furthermore, using (75) and (81), one can show that the relation (85) holds for arbitrary $k$.
@@ 4.5 Potential
$V_{n}(z)=z\left\langle{E_{n}\left|\frac{1}{1+z\bm{\rho}}\bm{\chi}\right|E_{0}}\right\rangle=z\left\langle{E_{0}\left|\frac{1}{1+z\bm{\rho}}\bm{\chi}\right|E_{n}}\right\rangle,$ (87)
@@ 4.6 Semiclassical approximation
At large $k$, the function $\psi(x|z)$ can be expanded perturbatively in powers of $1/k$, see (115) below. Each term in this expansion is polynomial in $x$ and $z$. This behaviour is consistent with the expected analytical properties of the function $\psi(x|z)$. Namely, for $k\to\infty$, the analyticity strip $\pi k<\mathop{\rm Im}\nolimits x<\pi k$ depicted in Figure 1 extends over the entire complex $x$-plane.
@@
A second complication arises from the structure of relation (3), which involves the function $\psi(x+2\pi in|z)$ evaluated for real $x$. For finite $k$ obeying the constraint (7), the point $x+2\pi in$ remains inside the analyticity strip, so the function $\psi(x+2\pi in|z)$ is well defined. In the semiclassical limit $k\to 0$, however, this point moves outside the strip. [^7] This issue is absent for $n=0$ and, in particular, does not affect the computation of the partition function. Consequently, the semiclassical evaluation of the generating function (3) requires an analytic continuation of the WKB solution (91) into the complex $x$-plane. Such a continuation is inherently ambiguous unless the analytic structure of $\psi(x|z)$ is fixed a priori. The computation of the Wilson loop in Klemm:2012ii adopted a specific prescription for this analytic continuation. The discrepancy with the numerical results, described in the Introduction, suggests that this choice is not correct.
@@ 5.2 Results
Combined with the relation (110), this leads to an exact expression for the generating function of the $1/2$ BPS Wilson loop in terms of the free energy $\mathcal{F}(z)$. We recall that the partition function $\Xi(z,k=4)$, and hence the free energy $\mathcal{F}(z)$, admits a closed-form representation in terms of Jacobi theta functions Grassi:2014uua. Together with relation (112), this yields an exact expression for the $1/2$ BPS Wilson loop at $k=4$.
@@ 5.3 Weak coupling regime
Substituting these expansions into (5.2) and (5.2) yields the large-$k$ behaviour of the coefficient functions $\mathcal{A}_{n}$, $\mathcal{B}_{n}$, and $\mathcal{C}_{n}$, which in turn allows us to determine the generating function (107).
@@ 6 Large $z$ regime
In this section, we determine the leading asymptotic behaviour of the generating function (107) in the limit $z\to\infty$, with the parameter $k$ held fixed. To analyse this regime, it is convenient to use the relation (11) and replace the fugacity parameter $z$ with the chemical potential $\mu$.
@@ 6.1 Coefficient functions at large $\mu$
$\mathcal{A}_{2}(z)=\frac{1}{4}e^{\frac{4\mu}{k}}+\dots\,$,
$\mathcal{B}_{2}(z)=-\frac{1}{4}e^{\frac{4\mu}{k}}\left[2\cot(\Phi(\mu)){\left({\cot({\textstyle\frac{2\pi}{k}})+\cot({\textstyle\frac{4\pi}{k}})}\right)}+1\right]+\dots\,$,
$\mathcal{C}_{2}(z)=-\frac{1}{2}e^{\frac{4\mu}{k}}{\left({\cot({\textstyle\frac{2\pi}{k}})+\cot({\textstyle\frac{4\pi}{k}})}\right)}+\dots\,.$ (123)
@@
The next step is to determine the large-$\mu$ behaviour of the derivative of the phase $\Phi^{\prime}(\mu)$.
@@ 6.2 Phase at large $\mu$
To determine the phase $\Phi(\mu)$, we complement relation (119) with the large-$\mu$ behaviour of the free energy $\mathcal{F}(z)$. Using the definitions (40) and (39), together with the expression (21) for the generating function $\Xi(z,k)$, we obtain
@@
$e^{\mathcal{F}(\mu)}{}$ $={e^{J^{\text{pert}}(\mu+2\pi im,k)}\over e^{J^{\text{pert}}(\mu-i\pi),k)}+e^{J^{\text{pert}}(\mu+i\pi),k)}}+O(e^{-{4\mu\over k}})$
@@
where, in the second step, we replaced $J^{\rm pert}(\mu)$ with its expression (23). Comparing this relation with (85) yields the large-$\mu$ behaviour of the phase,
@@ 7 Concluding remarks
In the above analysis, we have neglected the exponentially small corrections to (22) and (118) at large $N$. For large $k$ with fixed ’t Hooft coupling $\lambda=N/k$, in the type IIA string theory regime, the nonperturbative contribution (24) arises from string worldsheet instantons wrapping the $\mathbb{CP}^{1}$ inside $\mathbb{CP}^{3}$ Drukker:2010nc; Cagnazzo:2009zh, as well as from D2-brane instantons wrapping the $\mathbb{RP}^{3}$ cycle in $\mathbb{CP}^{3}$ Drukker:2011zy; Drukker:2010nc. In the M-theory regime, corresponding to large $N$ at fixed $k$, these two types of corrections uplift to distinct M2-brane instantons wrapping, respectively, $S^{3}/\mathbb{Z}_{k}$ (including the 11-dimensional circle) and $\mathbb{RP}^{3}\subset\mathbb{CP}^{3}$ inside $S^{7}/\mathbb{Z}_{k}$ (not including the 11-dimensional circle) Cagnazzo:2009zh; Giombi:2023vzu; Gautason:2023igo; Beccaria:2023ujc [^9] We are grateful to Arkady Tseytlin for useful discussions regarding this point.
@@
The nonperturbative corrections in ABJM theory were systematically investigated using the refined topological string representation in Nosaka:2015iiw; Hatsuda:2012dt; Hatsuda:2012hm; Hatsuda:2013oxa, and through direct computations of the contribution of quantum M2-branes in Giombi:2023vzu; Gautason:2023igo; Beccaria:2023ujc. High-precision numerical analyses in these works uncovered the intricate structure of instanton corrections to both the ABJM partition function Hatsuda:2015gca and supersymmetric Wilson loops Hatsuda:2013yua; Okuyama:2016deu. The method developed in this paper provides a way to understand this structure directly within the Fermi gas formalism, without relying on a conjectured duality with topological string theory BKT2.
@@
As mentioned above, the matrix model integral (5) obtained via localization is well-defined only for levels $k$ satisfying (7). For $k=2n$, the integral develops a pole. This pole does not appear in the weak-coupling expansion (5.3), but it arises for fixed $k$ due to the factor $1/\sin(2\pi n/k)$ in (118). In the large-$N$ limit with fixed $k$, the $1/2$ BPS Wilson loop admits a dual description in terms of an M2-brane wrapping the M-theory circle. It was shown in Giombi:2023vzu that for $n=1$ and $k>2$, the same factor $1/\sin(2\pi/k)$ is precisely reproduced by the one-loop contribution in the partition function of the wrapped M2-brane. Interestingly, for $k=1$ and $k=2$, the holographic prediction for the Wilson loop remains finite, while the localization result becomes singular. This discrepancy calls for an explanation.
@@
It would be interesting to generalize the method described in this paper to compute the latitude Wilson loop  [^10] The matrix model for the latitude Wilson loop was proposed in Bianchi:2018bke and tested at weak coupling up to three loops, both through Feynman-diagrammatic computations and via comparison with the bremsstrahlung function, showing nontrivial agreement. It was later derived in Griguolo:2021rke, where it was shown, under mild assumptions, that supersymmetric localization on $S^{3}$ reproduces the expected matrix integral. in ABJM theory Griguolo:2021rke; Bianchi:2018bke, as well as Wilson loops in the presence of deformations, such as real masses Armanini:2024kww; Nosaka:2015iiw and the squashing of the $S^{3}$ sphere, that break conformal invariance but preserve part of the supersymmetry. Notably, it has been conjectured (see Bobev:2025ltz and references therein) that the large $N$ partition function can still be expressed in terms of an Airy function, suggesting that Wilson loops may have the same property.
@@ Acknowledgements
We would like to thank Luca Griguolo, Luigi Guerrini, Shota Komatsu, Valentin Reys and Arkady Tseytlin for interesting discussions. We are indebted to Kazumi Okayama for sharing with us the results of his work Okuyama:2016deu.
The research of B.B. was supported by the Doctoral Excellence Fellowship Programme funded by the National Research Development and Innovation Fund of the Ministry of Culture and Innovation and the Budapest University of Technology and Economics, under a grant agreement with the National Research, Development and Innovation Offce (NKFIH). The work of G.K. and A.T. was supported by the French National Agency for Research grant “Observables” (ANR-24-CE31-7996).
@@ Appendix A Derivation of (3)
The relation (18) involves the matrix element (48), evaluated at $y=x+2\pi in$,
$\mathcal{W}_{n}(z,k)=\frac{\chi(x)}{4\pi k}\,\Gamma(x,x+2\pi in\,|\,z)\,.$ (131)
This matrix element admits an equivalent representation (49) in terms of the operator $\bm{K}$ defined in (50). The kernel of this operator satisfies the relation
$K(x,y)\,e^{y/k}+e^{x/k}K(x,y)=\frac{1}{4\pi k}\,E(x)E(y)\,$, (132)
where the function $E(x)$ is defined in (32). Multiplying both sides of this relation by $\chi(x)$, we can promote it to the operator identity
@@
where $\bm{\chi}=\chi(\bm{x})$ and $\langle x|\bm{\rho}|y\rangle=\chi(x)\,K(x,y)$. We can use this relation to get
@@ Appendix B Parity properties
where $\bm{\rho}^{t}=\bm{\chi}\bm{\rho}\bm{\chi}^{-1}$ denotes the transposed integral operator with kernel $\rho(y,x)$.
@@ Appendix C Properties of the potentials
The potentials $V_{nm}(z)$ are not independent. Taking matrix elements of both sides of (134) over the states $E_{n}$ and $E_{m}$, we obtain  [^11] Similar relations have previously appeared in Tracy_1996.

### v3

@@ 3.1 Properties of $\psi$-function
This operator possesses a discrete spectrum of real, positive eigenvalues that accumulate at the origin Hatsuda et al. (2012). These eigenvalues can be conveniently parameterized as $\lambda_{s}=e^{-\varepsilon_{s}}$, where $\varepsilon_{s}$ are monotonically increasing with $s$.
Using the operator representation (16), one finds that the eigenfunctions satisfy a finite-difference equation (see Marino (2017))  [^7] A conjecture for the exact eigenfunctions, based on the topological string/spectral theory correspondence, was proposed in Marino and Zakany (2017).
@@
with the expansion coefficients $c_{s}$ independent of $x$ and $z$. It then follows from the last relation that $\psi(x|z)$ is a meromorphic function of $z$ with poles located at the real positive $z=e^{\varepsilon_{s}}$.
@@
This function is entire in $z$ and has zeros located on the negative real axis, at the positions of the poles of the function $\psi(x|-z)$. For reasons that will become clear shortly, it is convenient to introduce the ratio of determinants
@@ 3.3 Wronskian relation
Inspection of the integral representation of the matrix elements on the right-hand side shows that the pole at $x-y=i\pi k$ arises solely from the $O(z)$ term. Matching the residues of (51) and (52) at this pole then yields the relation
@@ 4 Quantization condition
In the previous section, we expressed the generating function (3) in terms of the auxiliary $\psi$-function and derived a set of nontrivial relations satisfied by this function. These relations are exact and hold for arbitrary values of the Chern–Simons level $k$, the winding number $n$, and the fugacity parameter $z$.
@@ 4.3 Phase at odd $k$
$\begin{pmatrix}\psi(i\pi k|z)\\[5.69054pt] \psi(-i\pi k|z)\end{pmatrix}=\mathbb{T}_{\rm o}\begin{pmatrix}\psi(i\pi|z)\\[5.69054pt] \psi(-i\pi|z)\end{pmatrix}$. (78)
The corresponding transfer matrix reduces to $\mathbb{T}_{\rm o}=1$ for $k=1$, while for $k\geq 3$ it is given by
@@ 4.4 Solutions to the quantization conditions
It is straightforward to verify that the last two relations in (4.4) imply the same asymptotic behavior for $k=3,4$. Furthermore, using (75) and (81), one can show that the relation (85) holds for arbitrary $k$.
@@ 4.5 Potential
$V_{n}(z)=z\left\langle{E_{n}\left|\frac{1}{1+z\bm{\rho}}\bm{\chi}\right|E}\right\rangle=z\left\langle{E\left|\frac{1}{1+z\bm{\rho}}\bm{\chi}\right|E_{n}}\right\rangle,$ (87)
@@ 4.6 Semiclassical approximation
At large $k$, the function $\psi(x|z)$ can be expanded perturbatively in powers of $1/k$, see (115) below. Each term in this expansion is polynomial in $x$ and $z$. This behavior is consistent with the expected analytical properties of the function $\psi(x|z)$. Namely, for $k\to\infty$, the analyticity strip $\pi k<\mathop{\rm Im}\nolimits x<\pi k$ depicted in Figure 1 extends over the entire complex $x$-plane.
@@
A second complication arises from the structure of relation (3), which involves the function $\psi(x+2\pi in|z)$ evaluated for real $x$. For finite $k$ obeying the constraint (7), the point $x+2\pi in$ remains inside the analyticity strip, so the function $\psi(x+2\pi in|z)$ is well defined. In the semiclassical limit $k\to 0$, however, this point moves outside the strip. [^8] This issue is absent for $n=0$ and, in particular, does not affect the computation of the partition function. Consequently, the semiclassical evaluation of the generating function (3) requires an analytic continuation of the WKB solution (91) into the complex $x$-plane. Such a continuation is inherently ambiguous unless the analytic structure of $\psi(x|z)$ is fixed a priori. The computation of the Wilson loop in Klemm et al. (2013) adopted a specific prescription for this analytic continuation. The discrepancy with the numerical results, described in the Introduction, suggests that this choice is not correct.
@@ 5.2 Results
Combined with the relation (110), this leads to an exact expression for the generating function of the $1/2$ BPS Wilson loop in terms of the free energy $\mathcal{F}(z)$. We recall that the partition function $\Xi(z,k=4)$, and hence the free energy $\mathcal{F}(z)$, admits a closed-form representation in terms of Jacobi theta functions Grassi et al. (2016). Together with relation (112), this yields an exact expression for the $1/2$ BPS Wilson loop at $k=4$.
@@ 5.3 Weak coupling regime
Substituting these expansions into (5.2) and (5.2) yields the large-$k$ behavior of the coefficient functions $\mathcal{A}_{n}$, $\mathcal{B}_{n}$, and $\mathcal{C}_{n}$, which in turn allows us to determine the generating function (107).
@@ 6 Large $z$ regime
In this section, we determine the leading asymptotic behavior of the generating function (107) in the limit $z\to\infty$, with the parameter $k$ held fixed. To analyse this regime, it is convenient to use the relation (11) and replace the fugacity parameter $z$ with the chemical potential $\mu$.
@@ 6.1 Coefficient functions at large $\mu$
$\mathcal{A}_{2}=\frac{1}{4}e^{\frac{4\mu}{k}}+\dots\,$,
$\mathcal{B}_{2}=-\frac{1}{4}e^{\frac{4\mu}{k}}\left[2\cot(\Phi(\mu)){\left({\cot({\textstyle\frac{2\pi}{k}})+\cot({\textstyle\frac{4\pi}{k}})}\right)}+1\right]+\dots\,$,
$\mathcal{C}_{2}=-\frac{1}{2}e^{\frac{4\mu}{k}}{\left({\cot({\textstyle\frac{2\pi}{k}})+\cot({\textstyle\frac{4\pi}{k}})}\right)}+\dots\,.$ (123)
@@
The next step is to determine the large-$\mu$ behavior of the derivative of the phase $\Phi^{\prime}(\mu)$.
@@ 6.2 Phase at large $\mu$
To determine the phase $\Phi(\mu)$, we complement relation (119) with the large-$\mu$ behavior of the free energy $\mathcal{F}(z)$. Using the definitions (40) and (39), together with the expression (21) for the generating function $\Xi(z,k)$, we obtain
@@
$e^{\mathcal{F}(\mu)}{}$ $={e^{J^{\text{pert}}(\mu,k)}\over e^{J^{\text{pert}}(\mu-i\pi),k)}+e^{J^{\text{pert}}(\mu+i\pi),k)}}+O(e^{-{4\mu\over k}})$
@@
where, in the second step, we replaced $J^{\rm pert}(\mu)$ with its expression (23). Comparing this relation with (85) yields the large-$\mu$ behavior of the phase,
@@ 7 Concluding remarks
In the above analysis, we have neglected the exponentially small corrections to (22) and (118) at large $N$. For large $k$ with fixed ’t Hooft coupling $\lambda=N/k$, in the type IIA string theory regime, the nonperturbative contribution (24) arises from string worldsheet instantons wrapping the $\mathbb{CP}^{1}$ inside $\mathbb{CP}^{3}$ Drukker et al. (2011a); Cagnazzo et al. (2010), as well as from D2-brane instantons wrapping the $\mathbb{RP}^{3}$ cycle in $\mathbb{CP}^{3}$ Drukker et al. (2011b, a). In the M-theory regime, corresponding to large $N$ at fixed $k$, these two types of corrections uplift to distinct M2-brane instantons wrapping, respectively, $S^{3}/\mathbb{Z}_{k}$ (including the 11-dimensional circle) and $\mathbb{RP}^{3}\subset\mathbb{CP}^{3}$ inside $S^{7}/\mathbb{Z}_{k}$ (not including the 11-dimensional circle) Cagnazzo et al. (2010); Giombi and Tseytlin (2023); Gautason et al. (2023); Beccaria et al. (2023) [^10] We are grateful to Arkady Tseytlin for useful discussions regarding this point.
@@
The nonperturbative corrections in ABJM theory were systematically investigated using the refined topological string representation in Nosaka (2016); Hatsuda et al. (2013b, 2012, 2014), and through direct computations of the contribution of quantum M2-branes in Giombi and Tseytlin (2023); Gautason et al. (2023); Beccaria et al. (2023). High-precision numerical analyses in these works uncovered the intricate structure of instanton corrections to both the ABJM partition function Hatsuda et al. (2015) and supersymmetric Wilson loops Hatsuda et al. (2013a); Okuyama (2016a). The method developed in this paper provides a way to understand this structure directly within the Fermi gas formalism, without relying on a conjectured duality with topological string theory Boldis et al. (2026).
@@
As mentioned above, the matrix model integral (5) obtained via localization is well-defined only for levels $k$ satisfying (7). For $k=2n$, the integral develops a pole. This pole does not appear in the weak-coupling expansion (5.3), but it arises for fixed $k$ due to the factor $1/\sin(2\pi n/k)$ in (118). In the large-$N$ limit with fixed $k$, the $1/2$ BPS Wilson loop admits a dual description in terms of an M2-brane wrapping the M-theory circle. It was shown in Giombi and Tseytlin (2023) that for $n=1$ and $k>2$, the same factor $1/\sin(2\pi/k)$ is precisely reproduced by the one-loop contribution in the partition function of the wrapped M2-brane. Interestingly, for $k=1$ and $k=2$, the holographic prediction for the Wilson loop remains finite, while the localization result becomes singular. This discrepancy calls for an explanation.
@@
It would be interesting to generalize the method described in this paper to compute the latitude Wilson loop  [^11] The matrix model for the latitude Wilson loop was proposed in Bianchi et al. (2018b) and tested at weak coupling up to three loops, both through Feynman-diagrammatic computations and via comparison with the bremsstrahlung function, showing nontrivial agreement. It was later derived in Griguolo et al. (2021), where it was shown, under mild assumptions, that supersymmetric localization on $S^{3}$ reproduces the expected matrix integral. in ABJM theory Griguolo et al. (2021); Bianchi et al. (2018b), as well as Wilson loops in the presence of deformations, such as real masses Armanini et al. (2024); Nosaka (2016) and the squashing of the $S^{3}$ sphere, that break conformal invariance but preserve part of the supersymmetry. Notably, it has been conjectured (see Bobev et al. (2025) and references therein) that the large $N$ partition function can still be expressed in terms of an Airy function, suggesting that Wilson loops may have the same property.
@@ Acknowledgements
We would like to thank Luca Griguolo, Luigi Guerrini, Shota Komatsu, Valentin Reys and Arkady Tseytlin for interesting discussions. We are indebted to Kazumi Okayama for sharing with us the results of his work Okuyama (2016a). We are also grateful to Marcos Mariño for useful comments.
The research of B.B. was supported by the Doctoral Excellence Fellowship Programme funded by the National Research Development and Innovation Fund of the Ministry of Culture and Innovation and the Budapest University of Technology and Economics, under a grant agreement with the National Research, Development and Innovation Office (NKFIH). The work of G.K. and A.T. was supported by the French National Agency for Research grant “Observables” (ANR-24-CE31-7996).
@@ Appendix A Derivation of (3)
The relation (18) involves the function (48), evaluated at $y=x+2\pi in$,
$\mathcal{W}_{n}(z,k)=\int_{-\infty}^{\infty}dx\,e^{n(x+i\pi n)\over k}\frac{\chi(x)}{4\pi k}\,\Gamma(x,x+2\pi in\,|\,z)\,.$ (131)
This function admits an equivalent representation (49) in terms of the operator $\bm{K}$ defined in (50). The kernel of this operator satisfies the relation
$K(x,y)\,e^{y/k}+e^{x/k}K(x,y)=E(x)E(y)\,$, (132)
where the function $E(x)$ is defined in (32). Multiplying both sides of this relation by $\chi(x)/(4\pi k)$, we can promote it to the operator identity
@@
where $\bm{\chi}=\chi(\bm{x})$ and $\langle x|\bm{\rho}|y\rangle=\chi(x)\,K(x,y)/(4\pi k)$. We can use this relation to get
@@ Appendix B Parity properties
where $\bm{\rho}^{t}=\bm{\chi}^{-1}\bm{\rho}\bm{\chi}$ denotes the transposed integral operator with kernel $\rho(y,x)$.
@@ Appendix C Properties of the potentials
The potentials $V_{nm}(z)$ are not independent. Taking matrix elements of both sides of (134) over the states $E_{n}$ and $E_{m}$, we obtain  [^12] Similar relations have previously appeared in Tracy and Widom (1996a).